# Strategy for identifying Vera C. Rubin Observatory kilonova candidates for targeted gravitational-wave searches

### v1

@@ Abstract
Since the observation of the binary neutron star merger GW170817 and the associated kilonova AT2017gfo, the next joint gravitational-wave/optical kilonova has been highly anticipated. Overlapping observations between the Vera C. Rubin Observatory and the international gravitational-wave detector network are expected soon. Wide-field survey facilities, such as Rubin, can serve dual roles in gravitational-wave astronomy: conducting dedicated optical counterpart searches following gravitational-wave triggers and, through surveys such as the Legacy Survey of Space and Time (LSST), providing opportunities for fortuitous kilonova discoveries during routine operations. We use simulations to develop a strategy for identifying kilonova candidates observed by Rubin and processed by the Fink broker. These candidates can be used as astrophysical triggers for a targeted gravitational-wave search. We simulate kilonovae light-curves for the first year of Rubin with the latest observing strategy for the Wide-Fast-Deep and the Deep Drilling Fields. Assuming a kilonova rate of 250 Gpc-3 yr-1, we find that Rubin brokers should observe $\sim 4$ kilonovae per year with at least one alert above a signal-to-noise ratio of 5 within the gravitational-wave detector horizon ($\sim 350$ Mpc). Most of these will be faint, and detected 1–2 days following the neutron star merger. Photometric and spectroscopic follow-up will be limited to large telescopes. Using archival data from the Zwicky Transient Facility (ZTF) and our proposed selection criteria, we estimate a minimum contamination of at least 30 events per month from other transients and variables, even under our strictest selection criteria. A deep gravitational-wave search targeting Rubin kilonova candidates may lead to the next multi-messenger discovery.
@@
facilities: LIGO, Virgo, KAGRA, Rubinsoftware: This work made use of the following software packages: astropy (Astropy Collaboration et al., 2013, 2018, 2022), Jupyter (F. Perez & B. E. Granger, 2007; T. Kluyver et al., 2016), matplotlib (J. D. Hunter, 2007), numpy (C. R. Harris et al., 2020), python (G. Van Rossum & F. L. Drake, 2009), and scipy (P. Virtanen et al., 2020) and Fink (A. Möller et al., 2021; B. Biswas et al., 2023).
@@ 1 Introduction
The detection of gravitational waves (GWs) from the binary neutron star merger GW170817 (B. P. Abbott et al., 2017a) was a landmark discovery in astronomy. As well as being detected in gravitational waves, the event was detected across the electromagnetic spectrum, from radio to gamma-rays, with the short gamma-ray burst GRB 170817a detected $\sim 2$ s following the merger (B. P. Abbott et al., 2017b, c). The precise sky localisation allowed for the discovery of an optical counterpart, AT2017gfo, associated with the kilonova (B. P. Abbott et al., 2017b; D. A. Coulter et al., 2017; S. J. Smartt et al., 2017; I. Arcavi et al., 2017; I. Andreoni et al., 2017; M. Soares-Santos et al., 2017; N. R. Tanvir et al., 2017), allowing for GW170817 to be localised to its host galaxy, NGC 4993, at a distance of only 40 Mpc (B. P. Abbott et al., 2017b).
A second binary neutron star merger, GW190425 (B. P. Abbott et al., 2020a), was detected during the third observing run (O3; R. Abbott et al., 2021a, 2023a) of the International Gravitational Wave Detector Network (IGWN), consisting of Advanced LIGO (J. Aasi et al., 2015), Virgo (F. Acernese et al., 2015) and KAGRA (T. Akutsu et al., 2019, 2021). However, due to the poor sky localisation and larger distance compared to GW170817, no associated counterparts were found. The fourth LIGO-Virgo-KAGRA observing run is currently ongoing. Initial projections suggested that up to $10^{+52}_{-10}$ binary neutron stars may be detected (B. P. Abbott et al. 2020b, see also P. Petrov et al. 2022 and I. Andreoni et al., 2022a), but no significant binary neutron star candidates have been detected in low-latency in O4 to date111see https://gracedb.ligo.org/superevents/public/O4/ for gravitational-wave alerts from O4. W. Niu et al. (2025) present a potential sub-threshold binary neutron star merger in O4 (see also A. G. Abac et al., 2025a), though it was not discovered until almost two years after the gravitational-wave data were taken, which makes multi-messenger follow-up difficult. Detecting additional binary neutron star mergers with associated kilonovae [^2] In this paper, we restrict the discussion to binary neutron star mergers as progenitors to kilonovae. In principle, some fraction of neutron star black hole (NSBH) mergers may also produce gamma-ray bursts and kilonovae. However, mass ratios close to unity and high black hole spins are required to tidally disrupt the neutron star. Predictions based on the current constraints predict that only a small fraction of NSBH will tidally disrupt (e.g., S. Biscoveanu et al., 2023), and so we choose to ignore this class throughout this paper. Much of the discussion would generalise trivially to also include NSBH binaries. will allow for more precise constraints on cosmology, including measuring the Hubble constant (B. P. Abbott et al., 2017d; R. Abbott et al., 2023b; A. G. Abac et al., 2025b).
@@
Many strategies have been proposed for ToO observations to follow up binary neutron star merger candidates based on low-latency alerts from gravitational-wave detectors (e.g., S. Antier et al., 2020; I. Andreoni et al., 2022a; T. Ahumada et al., 2024). A common issue for follow up to gravitational-wave alerts is the large sky localisations of typical gravitational-wave events (B. P. Abbott et al., 2020b; P. Petrov et al., 2022) which are difficult to cover with small field of view instruments, along with the large distances (compared to GW170817), making the prospects of detecting faint kilonovae even more challenging (though see J. C. Rastinejad et al., 2022 and A. J. Levan et al., 2024 for kilonova candidates associated with gamma-ray bursts within 350 Mpc). Alternate methods, such as an optical search for orphan kilonovae, have also been explored (e.g., Z. Doctor et al., 2017; D. Scolnic et al., 2018; C. N. Setzer et al., 2019; I. Andreoni et al., 2019, 2020; V. Aivazyan et al., 2022; I. Tosta e Melo et al., 2024; N. Van Bemmel et al., 2025; M. D. Fulton et al., 2025).
@@
The Vera C. Rubin Observatory will soon enter full science operations with the Legacy Survey of Space and Time (LSST; Ž. Ivezić et al., 2019). Rubin is scheduled to begin science operations in October, 2025, coinciding with the final few months of the fourth observing run of LIGO-Virgo-KAGRA (O4; B. P. Abbott et al., 2020b). Rubin will image the Southern sky with a large field of view (almost 10 deg2) and is capable of imaging the sky to a limiting magnitude of $r<24.5$ with a $30$ s exposure. As part of LSST, it will survey the visible Southern sky every few days, revolutionising transient astronomy. However, this survey strategy is not ideal for rapidly evolving transients such as kilonovae (cf. I. Andreoni et al., 2022b), which have a peak brightness of $M_{AB}=-15$ magnitude and decline by up to a magnitude per night in $r$ band (and even faster in the bluer bands, assuming similar intrinsic behaviour to AT2017gfo), leading to only a few detections before the kilonova fades, making classification challenging. In addition, kilonovae are rare, and Rubin will produce millions of detections per night, consisting of supernovae, luminous red novae (G. Howitt et al., 2020), flare stars, variable sources, asteroids and kilonovae. This means there will be significant contamination for kilonovae searches (e.g., P. S. Cowperthwaite et al., 2018; M. D. Fulton et al., 2025; T. Barna et al., 2025). Due to the large volume of data being produced by Rubin, bespoke brokers such as Fink (A. Möller et al., 2021) have been developed, providing a way to process the data and classify transients.
@@
Gravitational waves from nearby binary neutron star mergers (within the current binary neutron star range; see Section 2.1) will be detected by low-latency automated search pipelines, as was the case for GW170817 (B. P. Abbott et al., 2017a). More distant binary neutron star mergers, with lower signal-to-noise ratios, may not be detected in online searches, but produce triggers in more sensitive offline searches (e.g., R. Magee et al., 2019; R. Abbott et al., 2024).
@@
In this paper, we develop a strategy for identifying kilonova candidates observed by Rubin and processed by Fink. These candidates can be used as astrophysical triggers to perform a targeted gravitational-wave search in LIGO/Virgo/KAGRA data. This work extends initial estimates done with early Rubin LSST observing strategies (C. N. Setzer et al., 2019) and includes difference image processing effects. By utilising the known location and approximately known time of a kilonova candidate, a targeted search for gravitational waves can be more sensitive than a standard all-sky search (cf. C. S. Kochanek & T. Piran, 1993; L. Z. Kelley et al., 2013), boosting the detection prospects for a joint gravitational-wave kilonova candidate by 40–100% (L. Z. Kelley et al., 2013; A. R. Williamson et al., 2014). This opens up the possibility of detecting sub-threshold binary neutron star mergers that may otherwise be missed in all-sky searches. In order to conduct such a search, a list of kilonova candidates needs to be compiled. We explore strategies for selecting kilonova candidates, estimate the number and properties of kilonovae detected by Fink/Rubin and estimate the level of contamination expected in such a sample.
@@ 2 Searches for gravitational waves from binary neutron star mergers
The LIGO-Virgo-KAGRA collaborations perform searches for gravitational waves from compact binary mergers in both a low-latency mode (with a goal of identifying a gravitational-wave transient as early as possible following merger, or even potentially prior to the merger; R. Magee et al. 2021) and a higher latency offline mode, which is typically more sensitive (e.g., A. G. Abac et al., 2025c, and references therein). These searches do not require any external trigger, and search all of the sky, all of the time when data from gravitational-wave detectors is available.
Targeted searches are also performed for a broad range of external astrophysical triggers from other messengers that may be associated with the emission of gravitational waves, such as gamma-ray bursts (R. Abbott et al., 2021b, 2022) and fast radio bursts (R. Abbott et al., 2023c). These external messengers typically identify the time and location of a potential source of gravitational waves that can be used to restrict the parameter space that is searched for gravitational-wave signals. The reduced parameter space allows for a more sensitive search (L. Z. Kelley et al., 2013).
For example, A. R. Williamson et al. (2014) find that in a targeted search for gravitational waves coincident with well-localised gamma-ray bursts, the distance to which signals could be detected was increased by a factor of up to 25%. Since the detection rate scales with the cube of the distance to which signals can be detected, this potentially leads to a doubling of the detection rate. In this paper, we suggest adopting a similar strategy to target kilonova candidates identified in data from the Rubin Observatory.
@@ 2.1 Binary neutron star merger rate and sensitive volume
Binary neutron stars mergers with a given set of parameters (masses, spins, location etc) produce a strain in a gravitational-wave detector. The signal-to-noise ratio ($\rho$) measures the amplitude of the signal strain relative to the noise in the detector. We assume that signals are detected if $\rho>\rho_{\mathrm{thresh}}$, assuming a single detector threshold of $\rho_{\mathrm{thresh}}=8$ (J. Abadie et al., 2010; S. Stevenson et al., 2015). The maximum distance that an optimally oriented and positioned binary neutron star merger is detectable to with $\rho=8$ is known as the horizon distance.
The binary neutron star inspiral range is defined as the angle-averaged distance at which a canonical binary neutron star signal (with two non-spinning neutron stars of mass $1.4$ M) can be detected. During O3, the inspiral range for both LIGO detectors was approximately 100–140 Mpc (R. Abbott et al., 2021a, 2023a). In O4a, the range was increased to around 160 Mpc (A. G. Abac et al., 2025a). At the time of writing, during O4c, the range is around 160 Mpc [^3] https://gwosc.org/detector_status/O4c/. The range is a factor of approximately 2.26 smaller than the horizon distance (J. Abadie et al., 2010), giving a horizon distance of $\sim 350$ Mpc.
@@ 3.1 Rubin LSST
The Vera C. Rubin Observatory is located in Cerro Pachón in Chile. Over the next ten years it will perform the Legacy Survey of Space and Time (LSST) which is designed to enable four primary science drivers: a census of the Solar System, a detailed map of the Milky Way, systematic exploration of the time-domain sky, and stringent constraints on dark matter and dark energy (Ž. Ivezić et al., 2019). Rubin LSST will obtain imaging of the Southern Sky in six band-passes, ugrizyugrizy, and limiting magnitudes above $24$ mag with nominal $30$ s exposures.
@@
LSST will revolutionize time-domain astronomy by uncovering an unprecedented number of astrophysical transients, ushering in a new era of optical big data. Variable sources will be detected through a difference-imaging pipeline, with transient candidates publicly reported within 60 seconds via a real-time alert stream. This system is projected to generate on the order of $10$ million alerts per night (E. Bellm et al., 2020).
@@ 3.2 Kilonova light-curve simulations
In order to estimate the population of kilonovae that will be observed by Rubin/LSST, we perform a series of simulations of kilonovae lightcurves. Simulations were generated using SNANA (R. Kessler et al., 2009) within the PIPPIN framework (S. Hinton & D. Brout, 2020). Throughout this work, we use SNANA calibrated flux (FLUXCAL) defined from the magnitude with a fixed zeropoint given by: $mag=27.5-2.5\log_{10}$(FLUXCAL).
For our simulations we use two types of kilonova light-curve models: the Kasen model (D. Kasen et al., 2017) used also in the PLAsTiCC classification challenge (R. Kessler et al., 2019), and the Bulla model (M. Bulla, 2019) which was used in the ELAsTiCC classification challenge. Kasen models are time-series spectral energy distribution models for binary neutron star mergers, parametrized by ejecta mass, ejecta velocity, and lanthanide fraction. Bulla models are based on radiation-transport Monte Carlo code models, and the emission includes two components of the kilonova ejecta: a high-opacity, lanthanide-rich component and a low-opacity, lanthanide-free component. These models have been compared to the observed emission of AT 2017gfo. Other extensions of this project, outside the scope of this work, would be to incorporate additional models such as the ones in N. Sarin et al. (2024).
We generate one year of LSST light-curves in ugrizyugrizy for each kilonova model. We repeat the simulation 100 times, equivalent to producing 4,261 light-curves per model. We use the observation strategy baseline v5 Ocean [^5] https://community.lsst.org/t/announcing-new-ocean-ddf-survey-strategy-simulation/10162 which contains both WFD and DDF. The difference between version 4.3 and Ocean is only in the DDFs which results in minimal changes in our results. We co-add exposures within $30s$ in the simulation to mimic the alert stream of Rubin. We generate kilonova light-curves in a redshift range of $0.015<z<0.11$. Lightcurves are simulated from 30 days before the peak to 60 days after the peak (in the rest frame) to capture any variability in the lightcurve on long timescales. When simulating kilonova light-curves we assume a rate of $250$ Gpc-3 yr-1 with no redshift dependence (A. G. Abac et al., 2025d). We find that on average Rubin will detect $42\pm 4$ kilonovae per year with this strategy.

### v2

@@ Abstract
Since the observation of the binary neutron star merger GW170817 and the associated kilonova AT2017gfo, the next joint gravitational-wave/optical kilonova has been highly anticipated. Overlapping observations between the Vera C. Rubin Observatory and the international gravitational-wave detector network are expected soon. Wide-field survey facilities, such as Rubin, can serve dual roles in gravitational-wave astronomy: conducting dedicated optical counterpart searches following gravitational-wave triggers and, through surveys such as the Legacy Survey of Space and Time (LSST), providing opportunities for fortuitous kilonova discoveries during routine operations. We use simulations to develop a strategy for identifying kilonova candidates observed by Rubin and processed by the Fink broker. These candidates can be used as astrophysical triggers for a targeted gravitational-wave search. We simulate kilonovae light-curves for the first year of Rubin with the latest observing strategy for the Wide-Fast-Deep and the Deep Drilling Fields. Assuming a kilonova rate of 250 Gpc-3 yr-1, we find that Rubin should observe $\sim 40$ kilonovae per year within the gravitational-wave detector horizon ($\sim 500$ Mpc). However, only $\sim 4$ will be received by the brokers with a signal-to-noise ratio larger than 5. Most of these will be faint, and detected 1–2 days following the neutron star merger. Photometric and spectroscopic follow-up will be limited to large telescopes. Using archival data from the Zwicky Transient Facility (ZTF) and our proposed selection criteria, we estimate a minimum contamination of at least 30 events per month from other transients and variables, even under our strictest selection criteria. A deep gravitational-wave search targeting Rubin kilonova candidates may lead to the next multi-messenger discovery.
@@
facilities: LIGO, Virgo, KAGRA, Rubin, ZTFsoftware: This work made use of the following software packages: astropy (astropy:2013; astropy:2018; astropy:2022), Jupyter (2007CSE.....9c..21P; kluyver2016jupyter), matplotlib (Hunter:2007), numpy (numpy), python (python), and scipy (2020SciPy-NMeth) and Fink (Moller:2021MNRASFINK; Biswas:2023A&AFinkKilonova).
@@ 1 Introduction
The detection of gravitational waves (GWs) from the binary neutron star merger GW170817 (Abbott:2017PRLGW170817) was a landmark discovery in astronomy. As well as being detected in gravitational waves, the event was detected across the electromagnetic spectrum, from radio to gamma-rays, with the short gamma-ray burst GRB 170817a detected $\sim 2$ s following the merger (Abbott:2017ApJGW170817Multimessenger; Abbott:2017ApJGW170817GRB170817A). The precise sky localisation allowed for the discovery of an optical counterpart, AT2017gfo, associated with the kilonova (Abbott:2017ApJGW170817Multimessenger; Coulter:2017Science; Smartt:2017Nature; Arcavi:2017Nature; Andreoni:2017PASAGW170817Australian; Soares-Santos:2017ApJL; Tanvir:2017ApJL), allowing for GW170817 to be localised to its host galaxy, NGC 4993, at a distance of only 40 Mpc (Abbott:2017ApJGW170817Multimessenger).
A second binary neutron star merger, GW190425 (Abbott:2020ApJGW190425), was detected during the third observing run (O3; Abbott:2021PhRvXGWTC2; Abbott:2023PhRvXGWTC3) of the International Gravitational Wave Detector Network (IGWN), consisting of Advanced LIGO (Aasi:2015CQGAdvancedLIGO), Virgo (Acernese:2015CQGAdvancedVirgo) and KAGRA (Akutsu:2019NatAstKAGRA; KAGRA:2021PTEP). However, due to the poor sky localisation and larger distance compared to GW170817, no associated counterparts were found. The fourth LIGO-Virgo-KAGRA observing run is currently ongoing. Initial projections suggested that up to $10^{+52}_{-10}$ binary neutron stars may be detected (Abbott:2020LRRObservingScenarios, see also Petrov:2022ApJ, Andreoni:2022ApJS and Kiendrebeogo:2023ApJ), but no significant binary neutron star candidates have been detected in low-latency in O4 to date111see https://gracedb.ligo.org/superevents/public/O4/ for gravitational-wave alerts from O4. Niu:2025arXiv present a potential sub-threshold binary neutron star merger in O4 (see also Abac:2025GWTC-4), though it was not discovered until almost two years after the gravitational-wave data were taken, which makes multi-messenger follow-up difficult. Detecting additional binary neutron star mergers with associated kilonovae [^2] In this paper, we restrict the discussion to binary neutron star mergers as progenitors to kilonovae. In principle, some fraction of neutron star black hole (NSBH) mergers may also produce gamma-ray bursts and kilonovae. However, mass ratios close to unity and high black hole spins are required to tidally disrupt the neutron star. Predictions based on the current constraints predict that only a small fraction of NSBH will tidally disrupt (e.g., Biscoveanu:2023MNRAS), and so we choose to ignore this class throughout this paper. Much of the discussion would generalise trivially to also include NSBH binaries. will allow for more precise constraints on cosmology, including measuring the Hubble constant (Abbott:2017Nature; Abbott:2023ApJ; Abac:2025GWTC4Cosmology).
@@
Many strategies have been proposed for ToO observations to follow up binary neutron star merger candidates based on low-latency alerts from gravitational-wave detectors (e.g., Antier:2020MNRAS; Andreoni:2022ApJS; Ahumada:2024PASP). A common issue for follow up to gravitational-wave alerts is the large sky localisations of typical gravitational-wave events (Abbott:2020LRRObservingScenarios; Petrov:2022ApJ; Kiendrebeogo:2023ApJ) which are difficult to cover with small field of view instruments, along with the large distances (compared to GW170817), making the prospects of detecting faint kilonovae even more challenging (though see Rastinejad:2022Nature and Levan:2024Nature for kilonova candidates associated with gamma-ray bursts within 350 Mpc). Alternate methods, such as an optical search for orphan kilonovae, have also been explored (e.g., Doctor:2017; Scolnic:2018; Setzer:2019MNRAS; Andreoni:2019PASP; Andreoni:2020; Aivazyan:2022; Melo:2024; vanBemmel:2025MNRASKNTRAP; Fulton:2025MNRAS).
@@
The Vera C. Rubin Observatory will soon enter full science operations with the Legacy Survey of Space and Time (LSST; Ivezic:2019ApJ). Rubin is scheduled to begin science operations in October, 2025, coinciding with the final few months of the fourth observing run of LIGO-Virgo-KAGRA (O4; Abbott:2020LRRObservingScenarios). Rubin will image the Southern sky with a large field of view (almost 10 deg2) and is capable of imaging the sky to a limiting magnitude of $r<24.5$ with a $30$ s exposure. As part of LSST, it will survey the visible Southern sky every few days, revolutionising transient astronomy. However, this survey strategy is not ideal for rapidly evolving transients such as kilonovae (cf. Andreoni:2022ApJSRubinCadence; Andrade:2025PASP), which have a peak brightness of $M_{AB}=-15$ magnitude and decline by up to a magnitude per night in $r$ band (and even faster in the bluer bands, assuming similar intrinsic behaviour to AT2017gfo), leading to only a few detections before the kilonova fades, making classification challenging. In addition, kilonovae are rare, and Rubin will produce millions of detections per night, consisting of supernovae, luminous red novae (Howitt:2020MNRAS), flare stars, variable sources, asteroids and kilonovae. This means there will be significant contamination for kilonovae searches (e.g., Cowperthwaite:2018ApJ; Fulton:2025MNRAS; Barna:2025arXiv). Due to the large volume of data being produced by Rubin, bespoke brokers such as Fink (Moller:2021MNRASFINK) have been developed, providing a way to process the data and classify transients.
@@
Gravitational waves from nearby binary neutron star mergers (within the current binary neutron star range; see Section 2.1) will be detected by low-latency automated search pipelines (Chaudhary:2024PNAS), as was the case for GW170817 (Abbott:2017PRLGW170817). More distant binary neutron star mergers, with lower signal-to-noise ratios, may not be detected in online searches, but produce triggers in more sensitive offline searches (e.g., Magee:2019ApJL; Abbott:2024PRDGWTC2.1).
@@
In this paper, we develop a strategy for identifying kilonova candidates observed by Rubin and processed by Fink. These candidates can be used as astrophysical triggers to perform a targeted gravitational-wave search in LIGO/Virgo/KAGRA data. This work extends initial estimates done with early Rubin LSST observing strategies (Setzer:2019MNRAS) and includes difference image processing effects. By utilising the known location and approximately known time of a kilonova candidate, a targeted search for gravitational waves can be more sensitive than a standard all-sky search (cf. Kochanek:1993ApJL; Kelley:2013PRD), boosting the detection prospects for a joint gravitational-wave kilonova candidate by 40–100% (Kelley:2013PRD; Williamson:2014PRDPyGRB). This opens up the possibility of detecting sub-threshold binary neutron star mergers that may otherwise be missed in all-sky searches. In order to conduct such a search, a list of kilonova candidates needs to be compiled. We explore strategies for selecting kilonova candidates, estimate the number and properties of kilonovae detected by Fink/Rubin and estimate the level of contamination expected in such a sample.
@@ 2 Searches for gravitational waves from binary neutron star mergers
The LIGO-Virgo-KAGRA collaborations perform searches for gravitational waves from compact binary mergers in both a low-latency mode (with a goal of identifying a gravitational-wave transient as early as possible following merger, or even potentially prior to the merger; Magee:2021ApJL) and a higher latency offline mode, which is typically more sensitive (e.g., Abac:GWTC4Methods, and references therein). These searches do not require any external trigger, and search all of the sky, all of the time when data from gravitational-wave detectors is available.
Targeted searches are also performed for a broad range of external astrophysical triggers from other messengers that may be associated with the emission of gravitational waves, such as gamma-ray bursts (Abbott:2021ApJLVKGRBO3a; Abbott:2022ApJLVKGRBO3b) and fast radio bursts (Abbott:2023ApJLVKFRBO3a). These external messengers typically identify the time and location of a potential source of gravitational waves that can be used to restrict the parameter space that is searched for gravitational-wave signals. The reduced parameter space allows for a more sensitive search (Kelley:2013PRD).
For example, Williamson:2014PRDPyGRB find that in a targeted search for gravitational waves coincident with well-localised gamma-ray bursts, the distance to which signals could be detected was increased by a factor of up to 25%. Since the detection rate scales with the cube of the distance to which signals can be detected, this potentially leads to a doubling of the detection rate. In this paper, we suggest adopting a similar strategy to target kilonova candidates identified in data from the Rubin Observatory.
@@ 2.1 Binary neutron star merger rate and sensitive volume
Binary neutron stars mergers with a given set of parameters (masses, spins, location etc) produce a strain in a gravitational-wave detector. The signal-to-noise ratio ($\rho$) measures the amplitude of the signal strain relative to the noise in the detector. We assume that signals are detected if $\rho>\rho_{\mathrm{thresh}}$, assuming a single detector threshold of $\rho_{\mathrm{thresh}}=8$ (Abadie:2010CQGRates; Stevenson:2015ApJ). The maximum distance that an optimally oriented and positioned binary neutron star merger is detectable to with $\rho=8$ is known as the horizon distance.
The binary neutron star inspiral range is defined as the angle-averaged distance at which a canonical binary neutron star signal (with two non-spinning neutron stars of mass $1.4$ M) can be detected. During O3, the inspiral range for both LIGO detectors was approximately 100–140 Mpc (Abbott:2021PhRvXGWTC2; Abbott:2023PhRvXGWTC3). In O4a, the range was increased to around 160 Mpc (Abac:2025GWTC-4). At the time of writing, during O4c, the range is around 160 Mpc [^3] https://gwosc.org/detector_status/O4c/. The range is a factor of approximately 2.26 smaller than the horizon distance (Abadie:2010CQGRates), giving a horizon distance of $\sim 350$ Mpc.
@@ 3.1 Rubin LSST
The Vera C. Rubin Observatory is located in Cerro Pachón in Chile. Over the next ten years it will perform the Legacy Survey of Space and Time (LSST) which is designed to enable four primary science drivers: a census of the Solar System, a detailed map of the Milky Way, systematic exploration of the time-domain sky, and stringent constraints on dark matter and dark energy (Ivezic:2019ApJ). Rubin LSST will obtain imaging of the Southern Sky in six band-passes, ugrizyugrizy, and limiting magnitudes above $24$ mag with nominal $30$ s exposures.
@@
LSST will revolutionize time-domain astronomy by uncovering an unprecedented number of astrophysical transients, ushering in a new era of optical big data. Variable sources will be detected through a difference-imaging pipeline, with transient candidates publicly reported within 60 seconds via a real-time alert stream. This system is projected to generate on the order of $10$ million alerts per night (ldm612).
@@ 3.2 Kilonova light-curve simulations
In order to estimate the population of kilonovae that will be observed by Rubin/LSST, we perform a series of simulations of kilonovae lightcurves. Simulations were generated using SNANA (Kessler:2009) within the PIPPIN framework (Hinton:2020). Throughout this work, we use SNANA calibrated flux (FLUXCAL) defined from the magnitude with a fixed zeropoint given by: $mag=27.5-2.5\log_{10}$(FLUXCAL).
For our simulations we use two types of kilonova light-curve models: the Kasen model (Kasen:2017) used also in the PLAsTiCC classification challenge (Kessler:2019), and the Bulla model (Bulla:2019) which was used in the ELAsTiCC classification challenge. Kasen models are time-series spectral energy distribution models for binary neutron star mergers, parametrized by ejecta mass, ejecta velocity, and lanthanide fraction. Bulla models are based on radiation-transport Monte Carlo code models, and the emission includes two components of the kilonova ejecta: a high-opacity, lanthanide-rich component and a low-opacity, lanthanide-free component. These models have been compared to the observed emission of AT 2017gfo. Other extensions of this project, outside the scope of this work, would be to incorporate additional models such as the ones in Sarin:2024.
We generate one year of LSST light-curves in ugrizyugrizy for each kilonova model. We repeat the simulation 100 times, equivalent to producing 4,261 light-curves per model. We use the observation strategy baseline v5 Ocean [^5] https://community.lsst.org/t/announcing-new-ocean-ddf-survey-strategy-simulation/10162 which contains both WFD and DDF. The difference between version 4.3 and Ocean is only in the DDFs which results in minimal changes in our results. We co-add exposures within $30s$ in the simulation to mimic the alert stream of Rubin. We generate kilonova light-curves in a redshift range of $0.015<z<0.11$. Lightcurves are simulated from 30 days before the peak to 60 days after the peak (in the rest frame) to capture any variability in the lightcurve on long timescales. When simulating kilonova light-curves we assume a rate of $250$ Gpc-3 yr-1 with no redshift dependence (Abac:2025GWTC-4Populations). In Figure 1 we show examples of generated light-curves for different models at various redshifts. We find that on average Rubin will observe $42\pm 4$ kilonovae per year with this strategy. However, as shown in Section  4.2.1 only a small percentage will have detections with signal-to-noise larger than 3 or 5.
Figure 1: Examples of simulated kilonova light-curves using Rubin observing strategy v5.0. We show examples both for Kasen and Bulla models at different redshifts (z). Circles indicate SNR>3>3 and lower triangles SNR<1<1 measurements for difference image photometry. A black vertical line indicates the simulated kilonova peak.